OpenStudy (anonymous):

find factors of 17^100 + 32^201

OpenStudy (aravindg):

is it 32 or 34?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

i think the statement is incorect

OpenStudy (anonymous):

vat do u think

OpenStudy (aravindg):

me 2

OpenStudy (anonymous):

so if there is 34 then , what would be the answer

OpenStudy (aravindg):

we can factor out 17 then

OpenStudy (anonymous):

how

OpenStudy (aravindg):

17*2=34

OpenStudy (unklerhaukus):

\[17^{100} + 32^{201}\]\[=17^{(10^2)} + (2^5)^{201}\]

OpenStudy (anonymous):

i have already done this but vat should be the next step

OpenStudy (unklerhaukus):

can you tell be where you found this question @singhmmm?

OpenStudy (anonymous):

i have the self tutor book

OpenStudy (unklerhaukus):

what is the chapter/sub-chapter

OpenStudy (anonymous):

polynomials

OpenStudy (anonymous):

so what are the factors?

OpenStudy (theviper):

It can't be FACTORED :)

OpenStudy (theviper):

It should be 34 instead of 32 for FACTORING :)

OpenStudy (shubhamsrg):

The rough factorization though would be: 17^100 + 32^201 It is of the form a^2 + 32b^4 where a= 17^50 and b=32^50 => a^2 + 2(2b)^4 =>a^2 + 2c^2 where c = 4b^2 or c=2^2 * 2^500 = 2^502 now a^2 + 2c^2 = a^2 + ( (sqrt2)c )^2 + 2sqrt2 ac - 2sqrt2 ac = (a+ sqrt2 c)^2 - [sqrt( 2sqrt2 ac)]^2 now again applying a^2-c^2 identity.. hmm.. long and tedious though but hey, its factorized! :P

OpenStudy (anonymous):

@shubhamsrg But \(\sqrt{2}\) is not an integer, and usually it's implied that we only want integer factors in these types of problems. Finding real factors is sort of trivial.

OpenStudy (shubhamsrg):

hmm, what you say is right. I just couldn't find another legitimate solution.